OpenStudy (anonymous):

9m^4-n^2 (factor the expression)

OpenStudy (anonymous):

use this formula a^2 - b^2 = (a - b) (a + b) here we have ( (3m^2)^2 - n^2 ) = ------------------------? here a = 3m^2 b = n

OpenStudy (anonymous):

(3m^2 +n) (3m^2-n) ?? Thanks! @tgawade

OpenStudy (anonymous):

Correct !! :)